OpenStudy (anonymous):

Graph the function y = 2/x. Draw a chart to show at least four ordered pairs that are on the graph. Its not divide in a fraction 2/x

OpenStudy (anonymous):

Ill give metals and fan for best answer.

OpenStudy (anonymous):

y=2/x looks like this: http://www.wolframalpha.com/input/?i=y%3D2%2Fx

OpenStudy (anonymous):

Or you can use substitution to find out what the ordered pairs are. f(0) = 2/0 = does not exist f(1) = 2/1 = 2 f(2) = 2/2 = 1 f(3) = 2/3 = 2/3 f(4) = 2/4 = 1/2 Does this make sense?

OpenStudy (anonymous):

um. sorta.

OpenStudy (anonymous):

I found the x and y values. You can then plot them on the graph. Do you need help making the graph? That is very hard to do online, but it is simply looking at the x and y values and basing it off of the graph that I showed you.

OpenStudy (anonymous):

i have a nother problem i need help with